# Long Video Diffusion Generation with Segmented Cross-Attention and Content-Rich Video Data Curation

### v1

@@ Abstract
Figure 1: Presto is capable of generating long videos with rich content and long-range coherence.
@@ 1 Introduction
Unlike short clips, long videos require a balance between content diversity and long-range coherence, posing significant challenges for current video generation methods. To generate long videos with rich content, we recognize the importance of expanding the text input by incorporating multiple texts, as demonstrated in previous approaches [18, 4]. While combining videos generated from each individual text can enhance content diversity, it often leads to abrupt transitions between different scenarios. One alternative is to simultaneously incorporate multiple texts into the video generation model. This method provides the model with a broader range of textual inputs, enabling the generation of more content-rich videos, in contrast to the traditional approach of using a single text input, which limits the available information. By modeling the texts concurrently, this approach helps ensure long-term coherence in the generated videos, providing a seamless viewing experience.
@@
Existing long video generation methods overlook the importance of high-quality data [4, 8], leading to low consistency and content diversity in generated videos. To implement a model capable of generating content-rich and coherent long videos, a large-scale, high-quality video dataset is crucial. This dataset must include content-rich, long videos with long-range coherence, along with multiple distinct yet coherent textual descriptions for each video. To this end, we develop a systematic data curation pipeline to collect content-rich video-prompt pairs from public datasets, contributing to the creation of our LongTake-HD dataset. We filter 261k single-scene video clips from 8.9M publicly accessible videos as the pre-training set. Then, we apply an additional set of meticulous filtering steps to select the finest instances, resulting in a fine-tuning dataset of 47k clips. This refined dataset ensures that our method can generate high-quality, extended-duration, and content-rich videos.
@@
To tackle the challenge of long videos, we propose Presto, a novel method capable of generating 15-second videos with rich content and long-range coherence, as shown in Fig. 1. Instead of using a single long caption for each video, we divide the visual content into segments and generate progressive sub-captions that align with the unfolding storyline. Next, we modify the cross-attention mechanism in the Diffusion Transformer (DiT). To adapt the DiT model for long video generation, we refine the text embedding process and the cross-attention mechanism to effectively handle multiple progressive text conditions alongside temporal information. In particular, we introduce Segmented Cross-Attention (SCA), which divides the hidden states into segments along the temporal dimension and cross-attends each segment with its corresponding sub-caption. SCA introduces no additional parameters or modules and can be seamlessly integrated into existing DiT-based methods with minimal fine-tuning. We explore three distinct SCA strategies to manage the interaction between text embeddings and segmented latent features: Isolate Segmented Cross-Attention (ISCA), Sequential Segmented Cross-Attention (SSCA), and Overlap Segmented Cross-Attention (OSCA). Our experiments demonstrate that OSCA enhances the content richness and long-range coherence in the generated long videos.
@@
The key contributions of our work are outlined as follows:
@@ 2 Related Work
Multiple-Text-to-Video Generation (MT2V) [4, 18, 22] aligns with our work, as we also incorporate multiple text prompts. While MT2V aims to create videos from multiple inputs, our model uniquely adheres to the traditional T2V framework by requiring only a single user prompt. Additionally, MT2V methods such as TALC [4], typically combine multi-scene captions rigidly (e.g., “A panda is running in the park, sunny.” and “A golden retriever is running in the park, autumn.”), while our progressive caption method eliminates redundant descriptions across sub-captions and focuses on scenario transitions. Our multiple sub-captions can coalesce into a continuous narrative, improving the coherence of the video content.
Video Generation Datasets are crucial for pre-training high-quality video generation models. Existing text-video datasets [3] have made substantial progress in terms of dataset size, such as Panda-70M [5], HD-VILA [42], and HD-VG [35], which contain 70M, 100M, and 130M video clips respectively. Recent works such as OpenVid-1M [16] and FlintstonesHD [44] have attempted to construct small, yet higher-quality datasets. In contrast, we propose LongTake-HD, focusing on the finest quality videos with rich content and long-range scenario coherence, as well as multiple progressive sub-captions per video.
@@ 3 LongTake-HD Dataset
Data curation plays a crucial role in training our proposed model. To generate content-rich videos with long-range coherence, it is essential to have both high-quality video content and detailed, descriptive captions. Beginning with a dataset of 8.9 million publicly available raw videos, we filtered it down to single-scene video clips that exhibit diverse content, with a resolution of 720×1280, a minimum duration of 15 seconds, and high aesthetic quality. This process yielded 261k instances, each comprising a content-rich video paired with five coherent and progressively structured sub-captions, as well as an overall video caption. To further enable dataset stratification for different training stages, we utilized the full set of 261k instances during the pre-training phase and applied rigorous filtering criteria to extract the finest quality 47k instances for the fine-tuning set. We refer to this curated dataset as LongTake-HD. Further details and specific thresholds are shown in Sec. A and Tab. 4 of Appendix.
@@ 3.1 Collecting Content-Diverse Video Clips
Existing publicly available datasets, such as HD-VILA-100M [42], HD-VG-130M [35], Panda 70M [5], and WebVid-10M [3], offer a vast array of diverse and comprehensive video data that reflect the natural distribution of real-world content. However, these raw datasets often contain substantial amounts of noisy and low-quality material, lacking in careful curation for content quality and caption coherence. Inadequate data curation processes can lead to the inclusion of noisy, disjointed, or irrelevant video data, which negatively impacts the training of models, particularly for long-form videos.
@@
Aesthetics and motion contents. We employ the LAION Aesthetics Predictor [28] to evaluate the aesthetic quality of video frames and remove those with low aesthetics scores. Optical flow is calculated using Unimotion [41], and clips with higher flow scores are retained to ensure a substantial level of motion amplitude. Furthermore, we compute the coefficient of variation for all optical flow values within each clip, defined as the ratio of the standard deviation to the mean [7]. This standardized measure of dispersion allows us to assess the smoothness and consistency of motion dynamics, helping to avoid abrupt shifts in motion intensity.
@@ 3.2 Obtaining Coherent Video Captions
We apply captioning techniques to both images and videos. First, we conduct image-level diversity filtering based on the sampled keyframes. Next, we generate captions for each keyframe and perform semantic filtering on these captions. Finally, we leverage Large Language Models (LLMs) to create multiple progressive sub-captions that include camera motions.
Diversity filtering for keyframes. We employ a comprehensive approach to evaluate the diversity and coherence of images, using a combination of low-level and semantic metrics. Specifically, we apply the Peak Signal-to-Noise Ratio (PSNR) [11] for pixel-wise filtering, the Structural Similarity Index Measure (SSIM) [38] for structure-wise filtering, and the Perceptual Similarity (LPIPS) [46] for semantic-wise filtering. Additionally, some sampled frames may contain minimal information, such as black screens or blurry images. To address this, we use the image file size as a filtering criterion, since frames with low information content typically result in smaller file sizes when compressed in the PNG format [49, 6, 50, 10].
Semantic filtering from captions. We utilize Large Vision-Language Models (LVLMs) as caption generators to create detailed descriptions for both the entire video and its sampled frames. The captions for individual frames offer in-depth descriptions of the visual elements in each keyframe and represent the corresponding short video segments, while the video caption emphasizes the dynamics and transitions across the video, incorporating both spatial and temporal details. To generate embeddings for all keyframe captions, we employ MPNet [31] from SentenceTransformers [24, 23] and compute the cosine similarity [30] between each pair of captions. Additionally, we filter out negative captions, which occur when LVLMs fail to generate responses for frames that contain sensitive or abstract content.
@@
As in the example above, the first sub-caption describes the main subject, ant, and the environment, soil. When it comes to the second and third sub-captions, it continues the previous story and elaborates further on the transitions, e.g. traverses the surface of the mound, and a larger view of the immediate ground area. This narrative-style sub-caption annotation strategy helps enhance the long-range coherence of the generated videos, and distinguishes our approach from existing video datasets and MT2V methods.
For a given long video, we first divide the video clip into $N$ segments and generate independent captions for each segment using our captioning model. We also create an overall description of the entire video, capturing both spatial and temporal details. Next, we refine each sub-caption using a causal approach. We employ a Large Language Model (LLM) to adjust each sub-caption, taking into account all previous sub-captions as well as the overall video description, ensuring that each sub-caption represents a distinct episode within the broader storyline. Additionally, we explicitly incorporate camera motion into each sub-caption to enable fine-grained control over the camera. This strategy results in a set of coherent and progressively linked sub-captions for diffusion model training.
During inference, when given a short prompt from the user, we again use the LLM as a “director” to generate $N$ scripts with consistent and detailed descriptions. For captioning both the full video and selected frames, we use Aria [13] as our captioning model, and GPT-4o [1] as the LLM for refining the sub-captions. The detailed prompt templates for these two models are provided in LABEL:lst:aria and LABEL:lst:gpt of Appendix.
@@ 4.1 Overview
We provide a brief overview of latent diffusion models (LDMs) for text-to-video generation. LDMs extend traditional transformer models to handle the generative task and conduct the diffusion process in the latent space. First, a pre-trained autoencoder is utilized to compress the raw image or videos from pixel to latent space and a text encoder takes the text input and creates text embeddings. A diffusion transformer (DiT) takes the visual input with noise and performs a denoising process during training. Specifically, as shown in Fig. 2(a), the diffusion transformer consists of a stack of self-attention and cross-attention blocks, which capture the spatial and temporal dependencies within the video as long as text embeddings condition. During inference, the diffusion transformer starts from an instance sampled from random Gaussian noise and applies the diffusion process iteratively across multiple timesteps, refining the output at each step.
To adapt the DiT model for long video generation, we modify the text embedding process and the cross-attention mechanism to effectively incorporate multiple progressive text conditions with temporal information. Specifically, we split the latent features into segments along temporal dimensions in the cross-attention mechanism and study three different strategies to implement the interaction between text embedding and segmented latent features. The proposed Segmented Cross-Attention (SCA), especially the overlap variant, improves the content richness and long-range coherence in generated long videos by a large margin. We study this proposed strategy and different variants in the next subsection.
@@ 4.2 Segmented Cross-Attention
Inspired by window attention[14], which limits the scope of attention to local regions, we propose the Segmented Cross-Attention (SCA) method. This method splits the hidden states into temporal-local segments to better interact with the progressive sub-captions via cross-attention. For each group of progressive sub-captions, we separately encode $N$ sub-captions with text encoder, and thus obtain a group of text embeddings $\{c_{i}\}_{i=1}^{N}\in N\times\mathbb{R}^{L\times D}$. Given the hidden state $z$ with $T$ frames in temporal dimension, we also evenly split $z$ into $N$ non-overlapped segments $\{z_{i}\}_{i=1}^{N}$ along the time dimension, aligning in quantity with the group of sub-captions. Thus, segment $z_{i}$ encapsulates the frame information ranging from $i\times T/N$ to $(i+1)\times T/N$. The core idea for our SCA is to constrain each segment $z_{i}$ to only access its corresponding text condition.
We study three strategies for segmented attention computing including: 1) Isolate Segmented Cross-Attention (ISCA), 2) Sequential Segmented Cross-Attention (SSCA), and 3) Overlap Segmented Cross-Attention (OSCA), as shown in Fig. 2(b).
For Isolate Segmented Cross-Attention (ISCA) shown in the first row of Fig. 2(b), we treat it as the basic setting of our SCA, in which each segment $z_{i}$ only cross-attends to its corresponding text condition $c_{i}$ in cross-attention layers. Due to the lack of internal interactions among segments, ISCA tends to generate video with rich content but lacks long-range coherence.
Another strategy to improve temporal coherence is to enable long-range interactions between latent features and text embeddings, as shown in the second row of Fig. 2(b). We refer this variant to Sequential Segmented Cross-Attention (SSCA), which concatenates the latent segment $z_{i}$ sequentially with its latter segments $\{z_{j}\}_{j>i}^{N}$ and takes the average of all attention outputs in the final. Such a strategy greatly improves the long-range coherence. However, the content richness drops because the sequential interactive introduces more information to each segment and blends its content diversity.

### v2

@@ Abstract
Figure 1: Presto can generate long videos with rich content and long-range coherence.
@@ 1 Introduction
Unlike short clips, long videos require a balance between content diversity and long-range coherence, posing significant challenges for current video generation methods. To generate long videos with rich content, we recognize the importance of expanding the text input by incorporating multiple texts, as demonstrated in previous approaches [22, 4]. While combining videos generated from each text can enhance content diversity, it often leads to abrupt transitions between different scenarios. One alternative is to incorporate multiple texts into the video generation model simultaneously. This method provides the model with a broader range of textual inputs, enabling the generation of more content-rich videos, in contrast to the traditional approach of using a single text input, which limits the available information. This approach helps ensure long-term coherence in the generated videos by modeling the texts concurrently, providing a seamless viewing experience.
@@
Existing long video generation methods overlook the importance of high-quality data [4, 9], leading to low consistency and content diversity in generated videos. A large-scale, high-quality video dataset is crucial to obtaining a model that generates content-rich and coherent long videos. This dataset should include long videos with long-range coherence and complex dynamics, along with multiple distinct yet coherent textual descriptions for each video. To this end, we develop a systematic data curation pipeline to collect content-rich video-prompt pairs from public datasets, contributing to creating our LongTake-HD dataset. We filter 261k single-scene video clips from 8.9M publicly accessible videos as the pre-training set. Then, we apply additional meticulous filtering steps to select the finest instances, resulting in a fine-tuning dataset of 47k clips. This refined dataset ensures that our method can generate high-quality, extended-duration, and content-rich videos.
@@
To tackle the challenge of long videos, we propose Presto, a novel method capable of generating 15-second videos with rich content and long-range coherence, as shown in Fig. 1. Instead of using a single long caption for each video, we divide the visual content into segments and generate progressive sub-captions that align with the unfolding storyline. Next, we modify the cross-attention mechanism in the Diffusion Transformer (DiT). To adapt the DiT model for long video generation, we refine the text embedding process and the cross-attention mechanism to effectively handle multiple progressive text conditions alongside temporal information. In particular, we introduce Segmented Cross-Attention (SCA), which divides the hidden states into segments along the temporal dimension and cross-attends each segment with its corresponding sub-caption. SCA introduces no additional parameters or modules and can seamlessly integrate into existing DiT-based methods with minimal fine-tuning. We explore three distinct SCA strategies to manage the interaction between text embeddings and segmented latent features: Isolate Segmented Cross-Attention (ISCA), Sequential Segmented Cross-Attention (SSCA), and Overlap Segmented Cross-Attention (OSCA). Our experiments demonstrate that OSCA enhances the content richness and long-range coherence in the generated long videos.
@@
Our key contributions are outlined as follows:
@@ 2 Related Work
Multiple-Text-to-Video Generation (MT2V) [4, 22, 27, 38] aligns with our work, as we also incorporate multiple text prompts. While MT2V aims to create videos from multiple inputs, our model uniquely adheres to the traditional T2V framework by requiring only a single user prompt. Additionally, MT2V methods such as TALC [4], typically combine multi-scene captions rigidly (e.g., “A panda is running in the park, sunny.” and “A golden retriever is running in the park, autumn.”). Our progressive caption method eliminates redundant descriptions across sub-captions and focuses on scenario transitions. That means our multiple sub-captions can coalesce into a continuous narrative, improving the coherence of the video content.
Video Generation Datasets are crucial for pre-training high-quality video generation models. Existing text-video datasets [3] have made substantial progress in terms of dataset size, such as Panda-70M [6], HD-VILA [48], and HD-VG [41], which contain 70M, 100M, and 130M video clips respectively. Recent works such as OpenVid-1M [20] and FlintstonesHD [50] have attempted to construct small, yet higher-quality datasets. In contrast, we propose LongTake-HD, focusing on the finest quality videos with rich content, long-range scenario coherence, and multiple progressive sub-captions per video.
Time-Varying Text Prompts. Notably, works like Phenaki [38] and VideoPoet [14] also explore the idea of utilizing the time-varying text prompts or latent to generate long videos, aligning with our methodology in high-level. A key difference is that our method presents a comprehensive solution to the long video generation problem, encompassing the dataset, model, and interpolation techniques, while these works primarily focus on the model aspect.
@@ 3 LongTake-HD Dataset
   Dimensions Video Captions Videos
   Dataset Caption Sub-Captions Tokens Duration Aesthetics Diversity†‡ Coherence†‡ Quality†‡
   Panda 13.2 8.5s 4.62 2.55 2.38 2.18
   HD-VILA 32.5 13.4s 4.78 2.52 2.49 2.31
   Ours ✓ (5) 186.42 15.7s 5.21 3.02 3.44 2.80
Table 1: Comparisons between popular text-video datasets and our LongTake-HD on different dimensions. Unless specifically noted otherwise, data is calculated over the entire dataset using automated metrics. Our dataset leads in all dimensions. : These aspects are evaluated on 100 random samples. : These aspects are evaluated via human reviews on a four-point scale.
Data curation plays a crucial role in training our proposed model. Both high-quality video content and detailed descriptive captions are essential for generating content-rich videos with long-range coherence. Beginning with a dataset of 8.9 million publicly available raw videos, we filtered it down to single-scene video clips that exhibit diverse content, with a resolution of 720×1280, a minimum duration of 15 seconds, and high aesthetic quality. This process yielded 261k instances, each comprising a content-rich video paired with five coherent and progressively structured sub-captions, as well as an overall video caption. To further enable dataset stratification for different training stages, we utilized the full set of 261k instances during the pre-training phase and applied rigorous filtering criteria to extract the finest quality 47k instances for the fine-tuning set. We refer to this curated dataset as LongTake-HD. Further details are shown in Appendix A.
@@ 3.1 Collecting Content-Diverse Video Clips
Existing publicly available datasets, such as HD-VILA-100M [48], Panda 70M [6], and WebVid-10M [3], offer a vast array of diverse and comprehensive video data that reflect the natural distribution of real-world content. However, these raw datasets often contain substantial amounts of noisy and low-quality material, lacking in careful curation for content quality and caption coherence. Inadequate data curation processes can lead to the inclusion of noisy, disjointed, or irrelevant video data, which negatively impacts the training of models, particularly for long-form videos.
@@
Aesthetics and motion contents. We employ the LAION Aesthetics Predictor [33] to evaluate the aesthetic quality of video frames and remove those with low aesthetics scores. Optical flow is calculated using Unimotion [47], and clips with higher flow scores are retained to ensure a substantial level of motion amplitude. Furthermore, we compute the coefficient of variation for all optical-flow values within each clip, defined as the ratio of the standard deviation to the mean [8]. This standardized measure of dispersion allows us to assess the smoothness and consistency of motion dynamics, helping to avoid abrupt shifts in motion intensity.
@@ 3.2 Obtaining Coherent Video Captions
We apply captioning techniques to both images and videos. First, we conduct image-level diversity filtering based on the sampled keyframes. Next, we generate captions for each keyframe and perform semantic filtering on these captions. Finally, we leverage Large Language Models to create multiple progressive sub-captions that include camera motions.
Diversity filtering for keyframes. We employ a comprehensive approach to evaluate the diversity and coherence of images, using a combination of low-level and semantic metrics. Specifically, we apply the Peak Signal-to-Noise Ratio (PSNR) [12] for pixel-wise filtering, the Structural Similarity Index Measure (SSIM) [44] for structure-wise filtering, and the Perceptual Similarity (LPIPS) [52] for semantic-wise filtering. Additionally, some sampled frames may contain minimal information, such as black screens or blurry images. To address this, we use the image file size as a filtering criterion since frames with low information content typically result in smaller file sizes when compressed in the PNG format [55, 7, 56, 11].
Semantic filtering from captions. We utilize Large Vision-Language Models as caption generators to create detailed descriptions for both the entire video and its sampled frames. The captions for individual frames offer in-depth descriptions of the visual elements in each keyframe and represent the corresponding short video segments, while the video caption emphasizes the dynamics and transitions across the video, incorporating both spatial and temporal details. To generate embeddings for all keyframe captions, we employ MPNet [36] from SentenceTransformers [29, 28] and compute the cosine similarity [35] between each pair of captions. Additionally, we filter out negative captions, which occur when LVLMs fail to generate responses for frames that contain sensitive or abstract content.
@@
As in the example above, the first sub-caption describes the main subject, ant, and the environment, soil. When it comes to the second and third, it continues the previous story and elaborates further on the transitions, e.g. traverses the surface of the mound, and a larger view of the immediate ground area. This narrative-style sub-caption annotation strategy helps enhance the long-range coherence of the generated videos and distinguishes our approach from existing video datasets and MT2V methods.
For a given long video, we first divide the video clip into $N$ segments and generate independent captions for each segment using our captioning model. We also obtain an overall description of the entire video, capturing both spatial and temporal details. Next, we refine each sub-caption using a causal approach. We employ an LLM to adjust each sub-caption, considering all previous sub-captions together with the overall description, ensuring that each sub-caption represents a distinct episode within the broader storyline. Additionally, we explicitly incorporate camera motion to enable fine-grained control over the camera. This strategy results in a set of coherent and progressively linked sub-captions for diffusion model training.
During inference, when given a short prompt from the user, we also use the LLM as a “director" to generate $N$ scripts with consistent and detailed descriptions. For captioning over videos and images, we use Aria [15] as our captioner and GPT-4o [1] as the LLM for refining the sub-captions. The detailed prompt templates for these two models are provided in LABEL:lst:aria and LABEL:lst:gpt of the Appendix. We also use an experiment to show the benefits of progressive sub-captions that may enhance the semantics in generated contents in Appendix B and Tab. 6.
3.3 Comparisons between Video Datasets
We compare our dataset with some popular text-video datasets, including HD-VILA-100M [48] and Panda-70M [6]. We evaluate both video captions and videos to show the high quality of our LongTake-HD. Results and details are shown in Tab. 1.
For video captions, while existing datasets typically offer a single overall video caption, our dataset includes an additional set of five time-varying sub-captions. These sub-captions are a key distinguishing feature of our LongTake-HD compared to others. Furthermore, these sub-captions can be directly concatenated to form a longer, comprehensive caption, aligning with the format of most text-video datasets. We also calculated the average number of tokens per video. The caption length significantly outperforms other datasets, being about six times longer than HD-VILA and fourteen times longer than Panda. We anticipate this characteristic will particularly benefit Diffusion model training, as highly descriptive captions have been proved crucial for text fidelity and video quality [5].
For videos, we began by calculating the average video duration. Our dataset demonstrates a significant advantage in average video length, primarily attributed to the exclusion of videos shorter than 15 seconds. Furthermore, we conducted a detailed analysis of video quality. Recognizing the complexity of video quality assessment, we concentrated our investigation on four key aspects: Aesthetics, Diversity, Coherence, and Quality. We randomly sampled 100 videos from each dataset to quantify these aspects and calculated average scores. Aesthetics, a commonly employed metric in evaluating video dataset quality, was assessed automatically using the LAION Aesthetics Predictor [33]. The latter three metrics were chosen for their alignment with the criteria utilized in our qualitative analysis and user study (see Sec. 5.3 and Tab. 3). Consequently, we opted for human evaluation, with reviewers scoring each video on a four-point scale. Our dataset exhibits a substantial improvement across all four quality metrics compared to other datasets, thereby highlighting the superior quality of the videos.
@@ 4.1 Overview
We provide a brief overview of latent diffusion models (LDMs) for text-to-video generation. LDMs extend traditional transformer models to handle the generative task and conduct the diffusion process in the latent space. First, a pre-trained autoencoder is utilized to compress the raw image or videos from pixel to latent space, and a text encoder takes the text input and creates text embeddings. A diffusion transformer (DiT) takes the visual input with noise and performs a denoising process during training. Specifically, as shown in Fig. 2(a), the diffusion transformer consists of a stack of self-attention and cross-attention blocks, which capture the spatial and temporal dependencies within the video as long as text embeddings condition. During inference, the diffusion transformer starts from an instance sampled from random Gaussian noise and applies the diffusion process iteratively across multiple timesteps, refining the output at each step.
To adapt the DiT model for long video generation, we modify the text embedding process and the cross-attention mechanism to effectively incorporate multiple progressive text conditions with temporal information. Specifically, we split the latent features into segments along temporal dimensions in the cross-attention, and study three different strategies to implement the interaction between text embedding and segmented latent features. The proposed Segmented Cross-Attention (SCA), especially the overlap variant, improves content richness and long-range coherence in generated long videos by a large margin. We will study this strategy and its different variants in the next subsection.
@@ 4.2 Segmented Cross-Attention
Inspired by window attention[18], which limits the scope of attention to local regions, we propose the Segmented Cross-Attention (SCA) method. This method splits the hidden states into temporal-local segments to better interact with the progressive sub-captions via cross-attention. For each group of progressive sub-captions, we separately encode $N$ sub-captions with text encoder, and thus obtain a group of text embeddings $\{c_{i}\}_{i=1}^{N}\in N\times\mathbb{R}^{L\times D}$. Given the hidden state $z$ with $T$ frames in the temporal dimension, we also evenly split $z$ into $N$ non-overlapped segments $\{z_{i}\}_{i=1}^{N}$ along the time dimension, aligning in quantity with the group of sub-captions. Thus, segment $z_{i}$ encapsulates the frame information ranging from $i\times T/N$ to $(i+1)\times T/N$. The core idea of our SCA is to restrict each segment $z_{i}$ to access only its corresponding text condition.
We study three strategies for segmented attention computing, as shown in Fig. 2(b). For Isolate Segmented Cross-Attention (ISCA) shown in the first row of Fig. 2(b), we treat it as the basic setting of our SCA, in which each segment $z_{i}$ only cross-attends to its corresponding text condition $c_{i}$ in cross-attention layers. Due to the lack of internal interactions between segments, ISCA tends to generate video with rich content but lacks long-range coherence.
Another strategy is to enable long-range interactions between latent features and text embeddings, as shown in the second row of Fig. 2(b). We refer to this variant as Sequential Segmented Cross-Attention (SSCA), which concatenates the latent segment $z_{i}$ sequentially with its latter segments $\{z_{j}\}_{j>i}^{N}$ and takes the average of all attention outputs in the final. Such a strategy greatly improves the long-range coherence. However, the content richness drops because the sequential interactive introduces more information to each segment and blends its content diversity.